OpenStudy (anonymous):

Find the first three terms of f(n + 1) = f(n) + 8, if f(1) = 27.

OpenStudy (freckles):

so the first term is given as 27 (where n=1) Now we are going to replace n with 2 to figure out the 2nd term.

OpenStudy (freckles):

oops actually let's replace n with 1 first this will give us the f(2) since 1+1=2

OpenStudy (freckles):

so replace all the n's in f(n+1)=f(n)+8 with 1's